1. .If gross profits are $300, and cost of goods sold is $600, what is the net sales?

2.

If the current account balances are:

Cash = $1,000

Accounts receivable = $500

Accounts payable = $750

Common Equity = 2,000

Fixed Assets = 1,500

Calculate current assets.

3. . If earnings before taxes (EBT) are $136,000, net sales (all on credit) are $315,000,   dividends are 25,000 and net income is 90,000, what is the tax expense?
Answer #1
1) Net sales = Gross profit+COGS = 300+600 = $             900
2) Current assets = Cash+Accounts receivable = 1000+500 = $          1,500
3) Tax expense = EBT-Net income = 136000-90000 = $       46,000